2 boats A and B move in perpendicular direction to a buoy anchored at some point O. They travel with velocity 1.2v, where v is the stream velocity. A moves along the river and B moves perpendicular to it. After traversing an equal distance from O the 2 boats return. Find the ratio of the time taken by the 2 boats 

Grade:11

1 Answers

Arun
6 years ago
The answer to this question is n/((n)2-1)1/2 (n stands for η)Solution:Assume the distance they transveresed=dFor A the velocity when he is gong with the flow of the river is nv+v and when he is coming against it is nv-vso the time required is (d/nv+v)+(d/nv-v) = 2nd/((nv)2-v2)For B, the velocity in going and getting back is ((nv)2-v2)1/2Fo the time is 2d/((nv)2-v2)1/2The TA/Tb= (2nd/(nv2-v2))/(2d/((nv)2-v2)1/2)=n/((nv)2-v2)1/2=n/((n)2-1)1/2